OpenStudy (anonymous):

The domain of the function g(x)=5/(1−e^x) is (in interval notation)?

OpenStudy (anonymous):

all real numbers except where \[1-e^x=0\] \[e^x=1\] \[x=0\]

OpenStudy (anonymous):

oh interval notation \[(-\infty,0)\cup (0,\infty)\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw